OpenStudy (anonymous):

Convert to an improper faction: 12 1/7

OpenStudy (zarkon):

12 1/7 =\[\frac{7*12+1}{7}\]

OpenStudy (anonymous):

12 1/7 = ((12*7) +1)/7 = 85/7

OpenStudy (anonymous):

Thanks ....